the probability that a radish seed willgerminate is 0.7. A gardener plants seeds in batches of 6. find the mean and the standard devation for the number of seeds germinating for each batch
Answer by stanbon(75887)

the probability that a radish seed willgerminate is 0.7.
A gardener plants seeds in batches of 6.
Find the mean and the standard deviation for the
number of seeds germinating in each batch
----
Binomial:
mean = np = 6*0.7 = 4.3
std = sqrt(npq) = sqrt(4.3*0.3) = 1.1358
=======================
Cheers,
Stan H.
=======================